OpenStudy (anonymous):

Brenda throws a dart at this square-shaped target: Part A: Is the probability of hitting the black circle inside the target closer to 0 or 1? Explain your answer. (5 points) Part B: Is the probability of hitting the white portion of the target closer to 0 or 1? Explain your answer. (5 points)

OpenStudy (misty1212):

guess what we need?

OpenStudy (anonymous):

A square is shown with sides labeled 11. A shaded circle is shown in the center of the square. The diameter of the circle is 2.

OpenStudy (misty1212):

never seen you out of your genie suit though

OpenStudy (anonymous):

I can't post the pic. :(

OpenStudy (misty1212):

ok we can still to it

OpenStudy (anonymous):

Lol yeah that's true lol

OpenStudy (anonymous):

ok

OpenStudy (misty1212):



OpenStudy (misty1212):

like that?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok this is about areas

OpenStudy (anonymous):

Lol I forgot I could draw it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the large square its 11 x 11 which = to 121

OpenStudy (anonymous):

but im not sure how to find the area of the circle :(

OpenStudy (misty1212):

the area of the circle is \(2\pi r=2\pi\) which is approximately \(6.28\) the area of the square is \(11^2=121\)

OpenStudy (anonymous):

oh so then the area of the circle is 6.28 and for the square its 121?

OpenStudy (anonymous):

correct?

OpenStudy (misty1212):

the area in the square outside the circle is \[121-6.28\] which is much larger than the area of the circle

OpenStudy (anonymous):

oh yeah true

OpenStudy (misty1212):

so the probability of getting inside the circle is closer to zero the probability of getting outside is closer to 1

OpenStudy (anonymous):

? i dont understand wat u just said : S @misty1212

OpenStudy (anonymous):

oh oh ok thanks lol soooooo much ur a lifesaver!!!!!!!!!!

OpenStudy (misty1212):

i take it you got it right?

OpenStudy (anonymous):

Yeah I did :) Now i just have to type a lot and I'm done!! Thx :)

OpenStudy (misty1212):

the probability of hitting inside the circle is about \[\frac{6.28}{121}\]not a very big number

OpenStudy (misty1212):

just nod your head and blink

OpenStudy (anonymous):

:) lol

OpenStudy (anonymous):

http://imgarcade.com/1/i-dream-of-jeannie-blink-gif/

OpenStudy (misty1212):

lml

OpenStudy (anonymous):

:D